OpenStudy (anonymous):

Vincent is giving cakes to some children at a carnival. He has 2 strawberry cakes, 6 pineapple cakes, and 7 chocolate cakes. If Vincent selects a cake randomly without looking, what is the probability that he will give a pineapple cake to the first child and then a strawberry cake to the second child? 6 over 15 multiplied by 2 over 14 is equal to 12 over 210 6 over 15 plus 2 over 14 is equal to 114 over 210 6 over 15 multiplied by 2 over 15 is equal to 12 over 225 6 over 15 plus 2 over 15 is equal to 8 over 15

OpenStudy (calculusxy):

I think that your first sentence with " 6 over 15 multiplied by 2 over 14 is equal to 12 over 210" should be correct. After all, this is a dependent variable. \[\frac{ 6 }{ 15 } \times \frac{ 2 }{ 14 } = \frac{ 12 }{ 210 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can I ask you one more

OpenStudy (calculusxy):

Sure. I'll try :)

OpenStudy (anonymous):

Celine is playing a game at school fair carnival. There is a box that has white green blue and orange marble there is also 12 sided die labled 1 though 12 how many outcomes are in a sample space for pulling out of the box and rolling a dice

OpenStudy (anonymous):

Answers are 48, 32,16,8

OpenStudy (calculusxy):

I am honestly not sure, but just an educated guess that it can be 48 because 12 x 4 .

OpenStudy (anonymous):

oh ok thank you

OpenStudy (calculusxy):

Hope I helped you :)